# Atypical generic directions in Teichmüller space

### v1

@@ Abstract.
Using work of Chaika–Masur–Wolf and Durham–Zalloum, we construct the first example of a sublinearly-Morse Teichmüller geodesic ray with minimal non-uniquely ergodic vertical foliation.
@@ 1. Introduction
There is delicate interplay between the geometry of the Teichmüller metric on the Teichmüller space, $\mathrm{Teich}(S)$, of a finite-type surface $S$ and its Thurston compactification, $\mathbb{P}\mathcal{MF}(S)$, by projectivized measured foliations. The crux of the complexity is that Teichmüller geodesics are defined by deformations of singular flat structures on $S$, while $\mathbb{P}\mathcal{MF}(S)$ is defined via hyperbolic geometry. This leads to an incongruity in how the internal geometry of $\mathrm{Teich}(S)$ is asymptotically encoded in $\mathbb{P}\mathcal{MF}(S)$. Notably, Lenzhen [Len08] proved that geodesic rays can have limit sets in $\mathbb{P}\mathcal{MF}(S)$ larger than a point (see also [CMW19, LLR18, BLMR20]).
In this note, we produce a concrete example highlighting how this tension manifests in the context of random walks of the mapping class group $\mathcal{MCG}(S)$ on $\mathrm{Teich}(S)$. For our purposes, the main question here is whether generic directions—namely geodesic rays tracked by sample paths of random walks—can be completely encoded via properties of the internal geometry of $\mathrm{Teich}(S)$.
@@
More recently, Gekhtman–Qing–Rafi [GQR22] proved that these tracking rays are sublinearly Morse [QR22, QRT24], a weak hyperbolicity condition which is controlled by the internal geometry of $\mathrm{Teich}(S)$ by work of Durham–Zalloum [DZ22]. In [GQR22], the authors proved that the set of accumulation points in $\mathbb{P}\mathcal{MF}(S)$ of all sublinearly Morse Teichmüller geodesic rays has full measure with respect to any stationary measure associated to a (sufficiently nice) random walk of $\mathcal{MCG}(S)$ on $\mathrm{Teich}(S)$. In other words, generic directions in $\mathrm{Teich}(S)$ are sublinearly Morse, and almost every sublinearly Morse geodesic is generic.
Our main result shows that sublinear Morseness does not characterize genericity:
@@
We end with an interesting consequence of Theorem A. There is a natural way to associate to any Teichmüller geodesic ray its limit set in $\mathbb{P}\mathcal{MF}(S)$. Notably, Cordes [Cor17] proved that this map is a well-defined injection when restricted to the set of Morse geodesic rays. That is, every Morse geodesic ray determines a unique limit point in $\mathbb{P}\mathcal{MF}(S)$. On the other hand, by Remark 2 and [CMW19, Theorem 2.7], the limit set in $\mathbb{P}\mathcal{MF}(S)$ of the Teichmüller geodesic rays in our examples can be made to be more than a point, including possibly an interval.
Corollary B.
There exist sublinearly-Morse Teichmüller geodesic rays whose limit sets in $\mathbb{P}\mathcal{MF}(S)$ are more than a point.
@@
Acknowledgments
We would like to thank Yushan Jiang and Abdul Zalloum for useful discussions. We would particularly like to thank Jon Chaika for providing some useful statements which helped us get started working with [CMW19]. Durham was partially supported by NSF grant DMS-1906487. Wu was partially supported by Simons collaboration grant 850685.
@@ 2. Informal discussion of [CMW19]
This choice of skewed vertical direction determines a lattice in the plane $\mathbb{R}^{2}$. The idea then is to choose a pair of points in the lattice which will correspond to the next vertical and horizontal directions that we want. Flowing along the Teichmüller geodesic flow, we can arrange that the lengths of these new vertical and horizontal lines become comparable to $1$, while the original slit becomes very long.
Now we have a very long slit on the torus, with one endpoint at the corner point (all four corners are identified), while the other point is somewhere in the interior. The last claim follows from arranging the original vertical direction to be skewed at an irrational angle, and also the flow time also being irrational, so that the other endpoint is not at the corner. This is the point of Proposition 1.
With this setup, we can draw an arc from the interior endpoint to the corner, whose length is bounded by the area of the torus. This new slit now lifts to a curve which has bounded length in the corresponding metric on $S$. We repeat this process, flowing to the next pair of vertices, so that the second slit now has very long length with one endpoint at the corner and the other in the interior.
@@ 3. Our Teichmüller geodesic and its slit curves
In this section, we show that the construction from [CMW19]. We begin by setting up some notation.
@@ 3.1. Setup and notation
Let $\alpha:=[1,4,9,16,25,...]$ be the real number with given continued fraction expansion. Let $T$ denote the horizontally and vertically oriented square torus and $Y=\begin{pmatrix}1&-\alpha\\ 0&1\end{pmatrix}T$. We glue two isometric, identically oriented copies of $Y$ together along a slit with holonomy $(\sum_{j=1}^{\infty}2(p_{2j}-q_{2j}\alpha),0)$. Denote this flat surface by $X$. Let $g_{t}=\begin{pmatrix}e^{t}&0\\ 0&e^{-t}\end{pmatrix}$, then the action of $g_{t}$ on $X$ induces a Teichmüller geodesic $\gamma$.
Observe that the first return of the vertical flow to the horizontal base is a rotation by $\alpha$. Let $\frac{p_{k}}{q_{k}}=[1,4,9,\dots,k^{2}]$.
@@
By [CMW19, Theorem 2.3] (originally done by Veech [Vee69]), if $\alpha=[a_{i}]$ has a subsequence $n_{k}$ such that $\sum_{k=1}^{\infty}a_{n_{k}+1}^{-1}<\infty$, then the associated interval exchange on the surface associated with the ergodic geodesic is non-uniquely ergodic. In our case, since $\alpha=[1,4,9,16,25,...]$, it follows that $\sum_{k=1}^{\infty}a_{k}^{-1}=\sum_{k=1}^{\infty}k^{-2}=\pi^{2}/6<\infty$. ∎
@@ 3.2. The geodesic ray and its slit curves
In this subsection, we prove begin our analysis of the geodesic ray $\gamma$. Roughly speaking, we need a sequence of times $\{t_{k}\}$—the gap between which grows logarithmically—so that at each time there is an increasingly short curve (the slit), in the complement of which there are no short curves. In Section 4, we show that this sequence of slit curves diverges in the curve graph $\mathcal{C}(S)$ at a linear rate (in the index of the sequence).
Our first Proposition 1 provides the sequence of times corresponding to the slit curves on the geodesic ray we need, then Proposition 2 shows that the slit curves that are sufficiently far apart are filling, and Proposition 3 in the next subsection providing the combinatorial information we need about how the short curves as those times are organized in $\mathcal{C}(S)$.
@@
    The (hyperbolic) lengths of the splitting slits $\zeta_{k}$ at time $t_{k}$ goes to 00 as $k\to\infty$.
@@
Moreover, by $|t_{k}-t_{k+D}|\leq|t_{k}-t_{k+1}|+\dots+|t_{k+D-1}-t_{k+D}|$, it follows that
@@
To see that $\log k<\log t_{k}$, recall that $\frac{p_{k}}{q_{k}}$ converges to the continued fraction $\alpha=[1,4,9,16,...]$ (hence each $q_{i}>0$), and recall $t_{k}:=\log(q_{2k+1})$, so $q_{k}=k^{2}q_{k-1}+q_{k-2}>k^{2}q_{k-1}$, so $\frac{q_{k}}{q_{k-1}}>k^{2}$. Therefore, $q_{k}>\prod_{i=1}^{k}i^{2}$, and hence $t_{k}>\log(q_{k})>\sum_{=1}^{k}\log(i^{2})>k$. In particular, $\log(t_{k})>\log(k)$.
@@
Item (3): Recall the shear of torus $\alpha$ on the torus, thus on the surface $X$, is given by $\alpha=[a_{i}]$, where $a_{i}=i^{2}$. By [CMW19, Lemma 2.16], if we set $n_{k}=2k+1$,
then the flat length of each slit $\zeta_{k}$ at time $t_{k}$ satisfies
@@
(Also see the proof of Theorem 2.7 of [CMW19, Section 7.1].) Therefore, $|\zeta_{k}|\to 0,$ as $k\to\infty$. For each slit $\zeta_{k}$, consider a disk $D_{k}$ in the torus containing the $\zeta_{k}$, which is indeed a cylinder. Gluing along the slit of each cylinder gives an annulus. By Riemann Mapping Theorem, this cylinder is conformal to a flat cylinder $C_{k}$.
Now, since flat length $|\zeta_{k}|$ is approaching 00, it implies the the modulus of a cylinder $C_{k}$, say $m_{k}$, would go to infinity. So the extremal length $E(\zeta_{k})$, (by definition, is the reciprocal of the modulus of the biggest cylinder) goes to 00, as $k$ goes to infinity. Now by Maskit’s comparison result [Mas85], we have
@@
Therefore, the hyperbolic length of the slit $\zeta_{k}$, $H(\zeta_{k})$, goes to 00, as $k$ goes to infinity. ∎
@@
By Lemma 1, the vertical flow of the flat surface $X$ defined above is non uniquely ergodic. The flat structure of $X$ corresponds to the average of the two ergodic measures ($c=0$ in the language of [CMW19]). If we replace the flat structure with one which correspond to a weighted sum of the two ergodic measure, where weights are non equal but both positive ($c\in(-1,1)$ and $c\not=0$ in the notation of [CMW19]), then for this new flat surface $X_{c}$, the conclusions of Proposition 1 are still valid. In particular:
@@
  • Item (3) is due to the fact that the horizontal component of the slit for $X_{0}=X$ is the average of the horizontal component of the slit for $X_{c}$ and for $X_{-c}$, hence the latter is no more than twice of the slit of $X_{0}$. The vertical components of these slits are all the same. Now apply the same argument as in Proposition 1 to get estimate of the hyperbolic length. Also see the proof of Theorem 2.7 in Section 7 of [CMW19].
To set ourselves up for our combinatorial arguments Section 4, we need the following observation:
@@
As before, denote the geodesic ray by $\gamma$. By [CMW19, Lemma 2.16], when $k$ is large enough, at time $t_{k+3}$, the horizontal length of $\zeta_{k+3}$ $\asymp 1/a_{n_{k+3}+1}=1/a_{2k+8}$, while the vertical length of $\zeta_{k+3}$ $\asymp q_{n_{k+2}}/q_{n_{k+3}}=q_{2k+5}/q_{2k+7}$. So, at time $t_{k+1}$, the vertical length of $\zeta_{k+3}$ would be much larger than $1$, hence would cross both torus vertically many times. In other words, for any point on the geodesic flow of $X$ at time $t_{k+1}$, the shortest horizontal segment from this point to $\zeta_{k+3}$ has length bounded by $O(1)$. We call this length the “horizontal distance”. Hence at time $t_{k}$, the shortest horizontal distance from every point to $\zeta_{k+3}$ $\asymp q_{n_{k}}/q_{n_{k+1}}=O(1/k^{4})$. By similar argument, at time $t_{k}$ the shortest vertical distance from every point to $\zeta_{k-3}$ is bounded by $O(1/k^{4})$. Because the slits are geodesics and the flat surface is non-positively curved, they intersect minimally. If there is some non-trivial loop on the surface which is disjoint from both slits, it has to travel outside a small neighborhood of the slit. However, by the argument above, at time $t_{k}$, outside a small neighborhood of the slit both tori are cut by $\zeta_{k-3}$ and $\zeta_{k+3}$ into a grid of small rectangles of size $O(1/k^{4})$, so such a non-trivial loop is impossible. This shows that $\zeta_{k-3}$ and $\zeta_{k+3}$ fill the surface. ∎

### v2

@@ Abstract.
Motivated by geometrically capturing generic directions in Teichmüller space—that is, tracking rays for random walks of the mapping class group—we use work of Chaika–Masur–Wolf and Durham–Zalloum to construct the first examples of a sublinearly-Morse Teichmüller geodesic rays with minimal non-uniquely ergodic vertical foliations.
@@ 1. Introduction
There is a delicate interplay between the geometry of the Teichmüller metric on the Teichmüller space, $\mathrm{Teich}(S)$, of a finite-type surface $S$ and its Thurston compactification, $\mathbb{P}\mathcal{MF}(S)$, by projectivized measured foliations. The crux of the complexity is that Teichmüller geodesics are defined by deformations of singular flat structures on $S$, while $\mathbb{P}\mathcal{MF}(S)$ is defined via hyperbolic geometry. This leads to an incongruity in how the internal geometry of $\mathrm{Teich}(S)$ is asymptotically encoded in $\mathbb{P}\mathcal{MF}(S)$. Notably, Lenzhen [Len08] proved that geodesic rays can have limit sets in $\mathbb{P}\mathcal{MF}(S)$ larger than a point (see also [CMW19, LLR18, BLMR20]).
In this article, we produce a concrete example highlighting how this tension manifests in the context of random walks of the mapping class group $\mathcal{MCG}(S)$ on $\mathrm{Teich}(S)$. For our purposes, the main question here is the following:
Question 1.
Can generic directions in $\mathrm{Teich}(S)$—namely geodesic rays tracked by sample paths of random walks—be completely characterized via properties of the internal geometry of $\mathrm{Teich}(S)$?
@@
More recently, Gekhtman–Qing–Rafi [GQR22] proved that these tracking rays are sublinearly Morse [QR22, QRT24], a weak hyperbolicity condition which is controlled by the internal geometry of $\mathrm{Teich}(S)$ (see Durham–Zalloum [DZ22]). In [GQR22], the authors proved that the set of accumulation points in $\mathbb{P}\mathcal{MF}(S)$ of all sublinearly Morse Teichmüller geodesic rays has full measure with respect to any stationary measure associated to a (sufficiently nice) random walk of $\mathcal{MCG}(S)$ on $\mathrm{Teich}(S)$. In other words, generic directions in $\mathrm{Teich}(S)$ are sublinearly Morse, and almost every sublinearly Morse geodesic is generic.
@@
The following is an interesting consequence of Theorem A:
Corollary B.
There exist divergent Teichmüller geodesic rays with the same underlying topological vertical foliations which diverge at a sublinear rate.
Proof.
In our examples, the simplex of measures supported by the (non-uniquely ergodic) vertical foliation is an interval. By Remark 2, we can we can construct two distinct rays $\gamma,\gamma^{\prime}$ each limiting to different endpoints in $\mathbb{P}\mathcal{MF}(S)$, and hence forcing $\gamma,\gamma^{\prime}$ to diverge. These rays, however, are in the same sublinearly Morse class by [DZ22, Theorem A, part (1)], which states that the sublinearly Morse boundary of Teichmüller space admits an injection into the Gromov boundary of the curve graph. Hence $\gamma,\gamma^{\prime}$ diverge at a sublinear rate. ∎
@@
We end with an interesting consequence of Corollary B. There is a natural way to associate any Teichmüller geodesic ray with its limit set in $\mathbb{P}\mathcal{MF}(S)$. Notably, Cordes [Cor17] proved that this map is a well-defined injection when restricted to the set of Morse geodesic rays. That is, every Morse geodesic ray determines a unique limit point in $\mathbb{P}\mathcal{MF}(S)$. On the other hand, by Remark 2 and [CMW19, Theorem 2.7], the limit set in $\mathbb{P}\mathcal{MF}(S)$ of the Teichmüller geodesic rays in our examples can be made to be more than a point, including possibly an interval.
@@ 2. Informal discussion of [CMW19]
This choice of skewed vertical direction determines a lattice in the plane $\mathbb{R}^{2}$. The idea then is to choose a pair of points in the lattice that will correspond to the next vertical and horizontal directions that we want. Flowing along the Teichmüller geodesic flow, we can arrange that the lengths of these new vertical and horizontal lines become comparable to $1$, while the original slit becomes very long.
Now we have a very long slit on the torus, with one endpoint at the corner point (all four corners are identified), while the other point is somewhere in the interior. The last claim follows from arranging the original vertical direction to be skewed at an irrational angle, and also the flow time being irrational, so that the other endpoint is not at the corner. This is the point of Proposition 1.
With this setup, we can draw an arc from the interior endpoint to the corner, whose length is bounded by the area of the torus. This new slit now lifts to a curve which has bounded length in the corresponding metric on $S$. We repeat this process, flowing to the next pair of vertices, so that the second slit now has a very long length with one endpoint at the corner and the other in the interior.
@@ 3. Our Teichmüller geodesic and its slit curves
In this section, we extract some basic information about the slit curves arising in the construction from [CMW19]. We begin by setting up some notation.
@@ 3.1. Setup and notation
Let $\alpha:=[1,4,9,16,25,\dots,k^{2},\dots]$ denote the real number with given continued fraction expansion, here by $\alpha=[a_{1},a_{2},a_{3},\dots]$ we mean
$\alpha=\frac{1}{a_{1}+\frac{1}{a_{2}+\frac{1}{a_{3}+\frac{1}{\dots}}}}$
and let $\frac{p_{k}}{q_{k}}$ denote the $k^{th}$ convergent for $\alpha$ from that expansion, i.e.
$\frac{p_{k}}{q_{k}}=[1,4,9,\dots,k^{2}]=\frac{1}{1+\frac{1}{4+\frac{1}{9+\frac{1}{\dots+\frac{1}{k^{2}}}}}}$
@@
Following the notation of [CMW19, Section 2], consider a subsequence $\{\frac{p_{n_{k}}}{q_{n_{k}}}\}$ defined by setting $n_{k}=2k+1$. Now as in [CMW19, Section 2.1], let
$b=2\sum_{k=1}^{\infty}\left(q_{n_{k}}\alpha-\lfloor q_{n_{k}}\alpha\rfloor\right)=2\sum_{k=1}^{\infty}\left(q_{2k+1}\alpha-p_{2k+1}\right)$
Let $T$ denote the horizontally and vertically oriented square torus and
$Y=\begin{pmatrix}1&-\alpha\\ 0&1\end{pmatrix}T$
We glue two isometric, identically oriented copies of $Y$ together along a slit with holonomy $(b,0)$. Denote this flat surface by $X$. Let
$g_{t}=\begin{pmatrix}e^{t}&0\\ 0&e^{-t}\end{pmatrix}$
then the action of $g_{t}$ on $X$ induces a Teichmüller geodesic $\gamma$. Observe that the first return of the vertical flow to the horizontal base is a rotation by $\alpha$.
At the beginning of Section 2 of [CMW19], the authors introduced three assumptions on the number $\alpha$, which we will now verify
Lemma 1.
The following hold:
  1. (A)
    $\sum_{k=1}^{\infty}(a_{{n_{k}}+1})^{-1}<\infty$:
  2. (B)
    $a_{n_{k}}\to\infty$:
  3. (C)
    $q_{n_{k-1}}\log a_{n_{k}+1}=o(q_{n_{k}})=o(a_{n_{k}}q_{n_{k}-1}+q_{n_{k}-2})$.
Proof.
For item (A), observe that
$\sum_{k=1}^{\infty}(a_{{n_{k}}+1})^{-1}=\sum_{k=2}^{\infty}(2k)^{-2}<\infty.$
For item (B), observe that $a_{n_{k}}=(2k+1)^{2}$ which converges to $\infty$ as $k\to\infty$.
Finally, for item (C), Remark 1 provides that
$\frac{q_{n_{k-1}}\log a_{n_{k}+1}}{q_{n_{k}}}=\frac{q_{2k-1}\log(a_{2k+2})}{q_{2k+1}}=O\left(\frac{\log(k)}{k^{4}}\right)=o(1).$
@@
This is an immediate consequence of [CMW19, Theorem 2.3], which was originally proved by Veech in [Vee69], which says that if $\alpha$ satisfies Assumption (A) in Lemma 1, then the vertical flow as an interval exchange map is non-uniquely ergodic. And Assumption (A) has been verified in Lemma 1. ∎
@@ 3.2. The geodesic ray and its slit curves
In this subsection, we begin our analysis of the geodesic ray $\gamma$. Roughly speaking, we need a sequence of times $\{t_{k}\}$, so that the gap between $t_{k},t_{k+1}$ grows logarithmically and so that at each time $t_{k}$ there is an increasingly short slit curves $\zeta_{k}$, in the complement of each of which there are no short curves. In Section 4, we show that this sequence of slit curves diverges in the curve graph $\mathcal{C}(S)$ at a linear rate (in the index of the sequence).
Our first Proposition 1 provides the sequence of times corresponding to the slit curves on the geodesic ray we need. Proposition 2 then shows that the slit curves whose indices are sufficiently far apart are filling. These results setup Proposition 3 in Section 4, which explains how the slits curves spread out at a uniform rate in the curve graph $\mathcal{C}(S)$.
@@
    The (hyperbolic) lengths of the splitting slits $\zeta_{k}$ at time $t_{k}$ goes to 0 as $k\to\infty$.
@@
Moreover, by
$|t_{k}-t_{k+D}|\leq|t_{k}-t_{k+1}|+\dots+|t_{k+D-1}-t_{k+D}|$
it follows that
@@
To see that $\log k<\log t_{k}$, recall that $\frac{p_{k}}{q_{k}}$ converges to the continued fraction for $\alpha=[1,4,9,16,$...] (hence each $q_{i}>0$). Also recall that $t_{k}:=\log(q_{2k+1})$, so
$q_{k}=k^{2}q_{k-1}+q_{k-2}>k^{2}q_{k-1}$
and
$\frac{q_{k}}{q_{k-1}}>k^{2}$
Therefore, $q_{k}>\prod_{i=1}^{k}i^{2}$, and hence
$t_{k}>\log(q_{k})>\sum_{=1}^{k}\log(i^{2})>k$
In particular, $\log(t_{k})>\log(k)$.
@@
Item (3): By [CMW19, Lemma 2.16], if we set $n_{k}=2k+1$, then the flat length of each slit $\zeta_{k}$ at time $t_{k}$ satisfies
@@
(Also see the proof of Theorem 2.7 of [CMW19, Section 7.1].) Therefore, $|\zeta_{k}|\to 0,$ as $k\to\infty$. For each slit $\zeta_{k}$, consider a disk $D_{k}$ in the torus containing the $\zeta_{k}$, which is indeed a cylinder. Gluing along the slit of each cylinder gives an annulus. By the Riemann Mapping Theorem, this cylinder is conformal to a flat cylinder $C_{k}$.
Now, since the flat length $|\zeta_{k}|$ of $\zeta_{k}$ is approaching 0, the modulus of a cylinder $C_{k}$, say $m_{k}$, must be diverging to $\infty$ with $k$. So the extremal length $E(\zeta_{k})$—which is, by definition, the reciprocal of the modulus of the biggest cylinder—goes to 0, as $k$ goes to infinity. Now by Maskit’s comparison result [Mas85], we have
@@
Therefore, the hyperbolic length of the slit $\zeta_{k}$, $H(\zeta_{k})$, goes to 0, as $k$ goes to infinity. ∎
@@
By Lemma 2, the vertical flow of the flat surface $X$ defined above is non uniquely ergodic. The flat structure of $X$ corresponds to the average of the two ergodic measures ($c=0$ in the language of [CMW19]). If we replace the flat structure with one which corresponds to a weighted sum of the two ergodic measures, where weights are non-equal but both positive ($c\in(-1,1)$ and $c\not=0$ in the notation of [CMW19]), then for this new flat surface $X_{c}$, the conclusions of Proposition 1 are still valid. Because the proof of sublinearly Morseness in the later sections only depends on Proposition 1, we can pick the Teichmuller geodesic whose limit is an interval in $\mathbb{P}\mathcal{MF}$, and the interval depends on the parameter $c$ (See [CMW19, Theorem 2.7], which requires that $\alpha$ satisfies Assumptions (A), (B) and (C) in Lemma1). In particular:
@@
  • Item (3) is due to the fact that the horizontal component of the slit for $X_{0}=X$ is the average of the horizontal component of the slit for $X_{c}$ and for $X_{-c}$, hence the latter is no more than twice that of the slit of $X_{0}$. The vertical components of these slits are all the same. Now apply the same argument as in Proposition 1 to get an estimate of the hyperbolic length. Also see the proof of Theorem 2.7 in Section 7 of [CMW19].
To set ourselves up for our combinatorial arguments in Section 4, we need another proposition.
@@
As before, denote the geodesic ray by $\gamma$. By [CMW19, Lemma 2.16], when $k$ is large enough, at time $t_{k+3}$, the horizontal length of the slit $\zeta_{k+3}$
$h(\zeta_{k+3})\simeq 1/a_{n_{k+3}+1}=1/a_{2k+8}$
(where $\simeq$ means bounded both above and below up to a multiplicative constant depending only on the surface), while the vertical length of the slit $\zeta_{k+3}$
$v(\zeta_{k+3})\simeq q_{n_{k+2}}/q_{n_{k+3}}=q_{2k+5}/q_{2k+7}$
So, at time $t_{k+1}$, the vertical length of $\zeta_{k+3}$ would be much larger than $1$, hence would cross both torus vertically many times. In other words, for any point on the geodesic flow of $X$ at time $t_{k+1}$, the shortest horizontal segment from this point to $\zeta_{k+3}$ has length bounded by $O(1)$, We call this length the “horizontal distance”. See Figure 2.
$P$
Figure 2. The blue dashed line represents the $\zeta_{k+3}$, and the length of the green segment represents the horizontal distance from some point $P$ to $\zeta_{k+3}$.
Hence, at time $t_{k}$, the shortest horizontal distance from every point to
$\zeta_{k+3}\simeq q_{n_{k}}/q_{n_{k+1}}=O(1/k^{4})$
By a similar argument, at time $t_{k}$ the shortest vertical distance from every point to $\zeta_{k-3}$ is bounded by $O(1/k^{4})$. Because the slits are geodesics and the flat surface is non-positively curved, they intersect minimally. If there is some non-trivial loop on the surface that is disjoint from both slits, it has to travel outside a small neighborhood of the slit. However, by the argument above, at time $t_{k}$, outside a small neighborhood of the slit both tori are cut by $\zeta_{k-3}$ and $\zeta_{k+3}$ into a grid of small rectangles of size $O(1/k^{4})$, so such a non-trivial loop is impossible. This shows that $\zeta_{k-3}$ and $\zeta_{k+3}$ fill the surface. ∎